Suppose you bought supplies for a party. Three rolls of streamers and 15 party hats cost $30. Later, you bought 2 rolls of streamers and 4 party hats for $11. How much did each roll of streamers cost? How much did each party hat cost?
Answer by stanbon(75887)

Suppose you bought supplies for a party.
Three rolls of streamers and 15 party hats cost $30.
Later, you bought 2 rolls of streamers and 4 party hats for $11.
How much did each roll of streamers cost?
How much did each party hat cost?
----------------------------------------
Equations:
3s + 15h = 30
2s + 4h = 11
--------------------
Multiply thru 1st by 2 ; Multiply thru 2nd by 3.
----
6s + 30h = 60
6s + 12h = 33
-------------------
Subtract 2nd from 1st:
18h = 27
h = 3/2 = $1.50 (price of each hat)
---------------------
Substitute into 2s + 4h = 11 to solve for "s".
2s + 4*(3/2) = 11
2s + 6 = 11
2x = 5
x = $2.50 (price of one streamer)
===========================================
Cheers,
Stan H.